### v3

# Exploring Two-Phase Continual Instruction Fine-tuning for Multilingual Adaptation in Large Language Models

Abstract
A key challenge for Large Language Models (LLMs) is improving their Multilingual instruction-following ability over time without deteriorating their ability in languages they already excel at, typically English. In this paper, we study a two-phase Continual Fine-tuning (CFT) setup toward improving a model’s Multilingual adaptability. Concretely, we consider a two-phase CFT process in which an English-only end-to-end instruction fine-tuned LLM (Phase 1) is sequentially fine-tuned on a multilingual instruction dataset (Phase 2). Across Mistral-7B and LLaMA-3-8B and multiple dataset pairs, we show that instructional similarity between phases is critical: aligned datasets preserve or improve English while boosting multilingual ability, whereas misaligned datasets cause English degradation. We show that this degradation arises from representation shift during CFT, and that targeted mitigation strategies, including generative replay and heuristic-based layer freezing, reduce this shift and improve multilingual adaptation.
Exploring Two-Phase Continual Instruction Fine-tuning for Multilingual Adaptation in Large Language Models
Divyanshu Aggarwalthanks: Joint First Author1, Sankarshan Damlefootnotemark: 1, Navin Goyal2, Satya Lokam1, Sunayana Sitaram2 1Microsoft  2Microsoft Research India {divyanshu.aggarwal,t-sandamle,navingo,satya.lokam,sunayana.sitaram}@microsoft.com
1 Introduction
Figure 1: Comparing t-SNEs (van der Maaten and Hinton, 2008) of the hidden activations for Mistral-7B and LLaMA-3-8B during our two-phase Continual Fine-tuning (CFT) process. We prompt each model with examples from MTBench Zheng et al. (2024), and visualize the similarity between the mean hidden activations, for each model layer. For datasets that encode "similar" instructions (Alpaca & MultiAlpaca), English ability does not decline (e.g., 3% gain for IFEval). For non-similar datasets (Instruct & MultiAlpaca), English ability declines (e.g., 8% decline for IFEval). Here, Phase 2 model representations do not align with Phase 1’s; thus, suggesting greater model weight interference and a decline in English ability.
The widespread adoption of Large Language Models (LLMs) has led to a growing multilingual user base Shiyas (2023). However, ensuring strong performance across languages remains a fundamental challenge, with models consistently performing worse on low-resource languages spoken by millions of speakers worldwide (Ahuja et al., 2023, 2024a). A key limitation is that both labeled and unlabeled training data are predominantly available in English and a few high-resource languages, while resources for other languages, especially low-resource ones, are scarce Shaham et al. (2024).
Training large models from scratch is computationally expensive, making fine-tuning pre-trained LLMs the preferred approach for improving multilingual capabilities Lankford et al. (2023); Nguyen et al. (2023). A common fine-tuning strategy is to train LLMs on an instruction-following dataset that contains a mixture of languages. However, these datasets are often heavily skewed toward English and other high-resource languages, leading to a performance imbalance: models perform strongly in English but struggle with low-resource languages Dhamecha et al. (2021); Li et al. (2024a, b). Further, prior works show that fine-tuning on a dataset that only contains non-English languages can hurt the model’s performance on English due to catastrophic forgetting Wu et al. (2025); Behrouz et al. (2025), which is not desirable for most real-world scenarios due to the volume of English queries Ta (2023). Ideally, we want the same model to be proficient in both English and other languages to avoid the costs of maintaining multiple models. We refer to an LLM’s proficiency in English as its English Ability (EA), and its effectiveness across other languages its Multilingual Ability (MA). In this work, we aim to improve an LLM’s MA while maintaining or improving its EA.
1.1 Our Approach
To bridge the gap between EA and MA, we introduce a two-phase Continual Fine-tuning (CFT) setup. We fine-tune a pre-trained LLM on an English instruction dataset in Phase 1 and then fine-tune it on a similarly-sized Multilingual dataset in Phase 2. In Phase 1, we use Alpaca Taori et al. (2023) and OpenOrca Lian et al. (2023), and in Phase 2 we use MultiAlpaca Wei et al. (2023) and mOpenOrca (§4.1). Alpaca and OpenOrca provide high-quality English instruction data, while MultiAlpaca and mOpenOrca are their multilingual counterparts, ensuring consistency in instruction style across phases. To compare the efficacy of our two-phase CFT setup, we compare it with a straightforward single-phase setup where the LLM is fine-tuned on the mixture of both the instruction tuning datasets.
We focus on two open-source models, LLaMA-3-8B and Mistral-7B as base models for our experiments. We also use fine-tuned versions of them, LLaMA-3-8B-Instruct and Mistral-7B-Instruct, as off-the-shelf Phase 1 English fine-tuned models [^1] LLaMA-3-8B’s pre-training data was 5% multilingual, but LLaMA-3-8B-Instruct is primarily non-multilingual (Dubey et al., 2024).. We quantify a model’s English Ability (EA) based on its performance on four English datasets: (i) Two datasets that measure instruction following capabilities (i.e., IFEval Zhou et al. (2023) and Alpaca Eval Li et al. (2023)) and (ii) two that measure reasoning abilities (i.e., MMLU Hendrycks et al. (2021) and HellaSwag Zellers et al. (2019)). Likewise, we quantify a model’s Multilingual Ability (MA) based on its performance on (i) two question-answering tasks (i.e., MLQA Lewis et al. (2019) and XQuAD Artetxe et al. (2019)) and (ii) XLSUM Hasan et al. (2021), a summarization task.
1.2 Our Contributions
CFT Outperforms Mixture.
We first observe that models trained using our two-phase CFT setup perform better than the single-phase "dataset mixture" setup (Tables 12; §4.2). Moreover, our two-phase CFT setup overall results in a better model for all languages, including English, for the same number of training steps. The two-phase CFT pipeline also provides more flexibility than training on a mixture of datasets, with the possibility of extending our approach to multi-phase fine-tuning, especially when data from earlier phases might not be available.
Forgetting vs. Dataset Similarity.
As mentioned earlier, fine-tuning with multilingual datasets to enhance a model’s multilingual ability can lead to a decline in its English ability due to catastrophic forgetting (Mukhoti et al., 2023; Winata et al., 2023). We investigate the factors that may lead to such forgetting by computing the similarity of English and Multilingual Instruction Fine-tuning (IFT) datasets. We observe that when English and multilingual datasets have instructions that are not similar, there is a decline in the Phase 2 model’s performance in English. On the other hand, when Phase 1 and Phase 2 datasets encode similar instructions, the Phase 2 model’s performance in English improves (refer to Figure 1). To quantify the similarity of these phase-wise datasets, we introduce two metrics based on language-agnostic embeddings and model representations. We show that our quantification correlates with the decline in English ability (Tables 34; §4.3).
Mitigating Forgetting.
We study the efficacy of two tailored variants of existing CFT strategies to mitigate the decline in EA after Phase 2 fine-tuning, while boosting MA. The first strategy is distribution replay. Here, we look at generative replay, i.e., using instructions from a similar English counterpart of the Phase 2 dataset to generate replay data using the Phase 1 model. We also try english replay which acts as language replay by utilizing existing English parallel data from the Phase 2 distribution. The second strategy employs layer freezing. Our heuristic selects specific layers for freezing during Phase 2 fine-tuning based on the weight differences between the Base and Phase 1 models. We also explore Spectrum Hartford et al. (2024) as an alternative heuristic. We study the gains in EA and MA of these strategies compared to specific baselines (Table 5; §5). To the best of our knowledge, we are the first to explore the effectiveness of CFT on LLMs with multilingual instruction datasets.
2 Related Work
Continual Learning in LLMs Chen et al. (2026).
In general, continual learning in LLMs can be broadly categorized into (i) continual pre-training (CPT) and (ii) continual fine-tuning (CFT). In CPT, the LLMs are continuously pre-trained to adapt to new domains or tasks by continuously updating them with new data alongside the existing data Shi et al. (2024). CPT builds on the existing LLM’s knowledge and is more computationally efficient than retraining an LLM using the current and old pre-training data Gupta et al. (2023). CPT is employed when distributional shifts occur (i) over time Amba Hombaiah et al. (2021); Jang et al. (2022a, b), (ii) across languages Jin et al. (2022); Fujii et al. (2024); Blevins et al. (2024) or (iii) across domains Ke et al. (2023); Gong et al. (2022); Xie et al. (2023).
On the other hand, CFT involves training the LLM on successive downstream tasks with varying data distribution or time shifts Shi et al. (2024). CFT comprises fine-tuning for different tasks Carrión and Casacuberta (2022); Guan et al. (2025), instruction-tuning Cahyawijaya et al. (2023); Kang et al. (2025), model refinement/editing Zhang et al. (2023) and alignment Suhr and Artzi (2023). Recent literature also focuses on using CFT to assist the LLM to learn new languages Praharaj and Matveeva (2023); Pfeiffer et al. (2022); Badola et al. (2023); Singh et al. (2024c).
CFT: Enhancing LLMs Multilingual Abilities.
Cahyawijaya et al. Cahyawijaya et al. (2023) propose InstructAlign which uses cross-lingual alignment and episodic replay to align an LLM’s pre-trained languages to unseen languages but requires parallel data and previous task data. Shaham et al. Shaham et al. (2024) introduces multilinguality during the first instruction fine-tuning phase which improves an LLM’s instruction following capability across languages. He et al. He et al. (2023) show catastrophic forgetting during CFT and use techniques such as joint fine-tuning and model regularization to mitigate it. However, these techniques are computationally expensive or require access to previous task data.
Multilingual Adaptation.
This set of works looks at language and task adaption by adjusting the model to understand new languages and enhancing its performance on specific tasks through fine-tuning, respectively Chen et al. (2023); Zhao et al. (2024); Pfeiffer et al. (2020). For instance, Chen et al. (2023) perform task adaption by fine-tuning the model on downstream task data. For language adaption, they fine-tune only the token embedding layer, helping the model learn specific lexical meanings of new languages. Language and english ability are either trained in parallel or sequentially. However, in this paper, we try to incorporate multilingual ability in models with the constraint that they may have already learned english ability (e.g., Mistral-7B-Instruct). To the best of our knowledge, this is a first attempt at studying the effect of task and language self-instruct datasets on an LLM’s multilingual ability through CFT.
3 Two-phase Continual Fine-tuning Setup
When instruction fine-tuning LLMs, the most natural method is to fine-tune on a "dataset mixture" containing English and Multilingual data Workshop et al. (2023). However, fine-tuning on all languages simultaneously may introduce performance bias where the model performs better in English (and other high resource languages) Dhamecha et al. (2021); Li et al. (2024a, b) [^2] In §4.2, we compare dataset mixture to CFT..
Continual Fine-tuning (CFT).
To improve the multilingual performance of pre-trained LLMs, we introduce the following two-phase CFT process.
Two-Phase CFT Process Phase 1: Fine-tune a base LLM end-to-end on an English instruction dataset. Phase 1 aims to teach the LLM English Instruction Following Ability, which we refer to as English Ability (EA). Phase 2: Take the fine-tuned LLM from Phase 1 and further fine-tune it end-to-end on a Multilingual instruction dataset. Phase 2 focuses on enhancing the LLM’s Multilingual Ability (MA), using a dataset with multiple languages and fewer data points per language.
Challenges.
The primary challenge in our two-phase CFT process is that the LLM’s Multilingual Ability must not come at the cost of its English Ability. We impose two additional constraints based on real-world scenarios. First, in Phase 2, we cannot re-use Phase 1’s dataset. Often instruction fine-tuned LLMs are available without their corresponding datasets (e.g., Mistral-7B-Instruct Jiang et al. (2023)). Second, in Phase 2, we cannot use the weights of the Phase 1 model during training, as saving both old and new set of parameters on the GPU for training would be computationally expensive.
Two-phase Continual Fine-tuning Model Phase 1 (P1) Phase 2 (P2) IFEval $(\uparrow)$ Alpaca Eval $(\uparrow)$ MMLU $(\uparrow)$ HellaSwag $(\uparrow)$ XLSUM_en $(\uparrow)$ Average Dataset Dataset P1 P2 P1 P2 P1 P2 P1 P2 P1 P2 P1 P2 Mistral-7B Alpaca 0.364 0.395 0.12 0.16 0.552 0.573 0.581 0.616 0.10 0.11 0.343 0.371 Instruct Multi 0.550 0.462 0.35 0.15 0.575 0.533 0.641 0.416 0.13 0.10 0.449 0.332 LLaMA-3-8B Alpaca Alpaca 0.277 0.326 0.10 0.11 0.231 0.242 0.556 0.567 0.07 0.08 0.247 0.265 Instruct 0.735 0.182 0.14 0.10 0.340 0.239 0.533 0.278 0.11 0.09 0.372 0.178 Dataset Mixture Model Dataset Mixture IFEval $(\uparrow)$ Alpaca Eval $(\uparrow)$ MMLU $(\uparrow)$ HellaSwag $(\uparrow)$ XLSUM_en $(\uparrow)$ Average Mistral-7B Alpaca MultiAlpaca 0.394 0.23 0.538 0.602 0.09 0.371 LLaMA-3-8B Alpaca MultiAlpaca 0.363 0.07 0.598 0.602 0.04 0.335
Table 1: English Ability results for two-phase Continual Fine-tuning (CFT). When the phase-wise datasets are similar (Definition 1 and Definition 2), English Ability post Phase 2 (P2) fine-tuning consistently improves (denoted with green). When the phase-wise datasets are not similar, we see a significant decline in English Ability post Phase 2 (P2) fine-tuning (denote with red). We also provide numbers for dataset mixture – when the models are fine-tuned simultaneously on the Phase 1 and Phase 2 datasets.
4 Evaluating English & Multilingual Ability for Multilingual CFT
4.1 Experiment Setup & Evaluation Tasks
Fine-tuning Models.
We continually fine-tune open-source Mistral-7B Jiang et al. (2023) and LLaMA-3-8B Dubey et al. (2024) LLMs for multilingual adaptation.
Fine-tuning Datasets.
For our phase-wise datasets, we use the open-source Alpaca Taori et al. (2023), MultiAlpaca Wei et al. (2023), and OpenOrca Lian et al. (2023) datasets. Alpaca is a self-instruct English-only dataset. MultiAlpaca is a multilingual dataset created by translating Alpaca’s seed tasks to 11 languages and using GPT-3.5-Turbo for response collection. The languages are in equal proportions and are “French”, “Arabic”, “German”, “Spanish”, “Indonesian”, “Japanese”, “Korean”, “Portuguese”, “Russian”, “Thai”, and “Vietnamese”. The appendix (§A.2) describes OpenOrca and mOpenOrca.
Fine-tuning Technique.
We perform full fine-tuning in bf16 precision to study the impact of multilingual post-training in Phase 2 on English ability. This setting allows us to fully leverage model capacity, which may not be attainable under parameter-efficient fine-tuning methods Aggarwal et al. (2024); Panda et al. (2024). To mitigate the resulting degradation in English performance, §5 introduces a heuristic-based layer freezing strategy that selectively freezes a subset of layers while fine-tuning the remainder. All experiments are conducted using Axolotl333github.com/axolotl-ai-cloud/axolotl/, an open-source framework for large-scale LLM fine-tuning.
Evaluation Tasks.
To quantify an LLM’s english ability, we evaluate Phase 1 and Phase 2 models on two instruction-following tasks (i) IFEval Zhou et al. (2023) and (ii) Alpaca Eval Li et al. (2023), (iii) MMLU Hendrycks et al. (2021) for problem-solving, (iv) HellaSwag Zellers et al. (2019) for commonsense reasoning ability, and (v) XLSUM_en Hasan et al. (2021) for summarization. To quantify an LLM’s multilingual ability, we evaluate our fine-tuned models on three benchmark datasets comprising two multilingual generative tasks: question answering (MLQA Lewis et al. (2019) & XQuAD Artetxe et al. (2019)), instruction-following (GMMLU Singh et al. (2024a)), and summarization (XLSUM Hasan et al. (2021)).
Our evaluation suite spans diverse linguistic phenomena and reasoning skills to comprehensively assess language understanding, generation, and problem-solving. It also incorporates parallel benchmarks for English and multilingual ability, enabling direct cross-lingual comparison. Further details are provided in §A.3.
To evaluate our models on EA and MA, we use LM-Evaluation-Harness444github.com/EleutherAI/lm-evaluation-harness, which is a unified framework for zero/few-shot evaluations of LLMs. For both English and multilingual ability, we use zero-shot evaluation. For additional details on the training setup, code, and evaluation tasks, refer to §A.
4.2 Results
Two-phase Continual Fine-tuning Model Phase 1 (P1) Phase 2 (P2) MLQA $(\uparrow)$ XLSUM $(\uparrow)$ XQuAD $(\uparrow)$ GMMLU $(\uparrow)$ Average Dataset Dataset P1 P2 P1 P2 P1 P2 P1 P2 P1 P2 Mistral-7B Alpaca 0.229 0.288 0.012 0.060 0.290 0.602 0.42 0.40 0.238 0.338 Instruct Multi 0.246 0.307 0.012 0.033 0.351 0.436 0.44 0.43 0.262 0.302 LLaMA-3-8B Alpaca Alpaca 0.438 0.597 0.033 0.034 0.586 0.737 0.53 0.34 0.397 0.427 Instruct 0.609 0.321 0.048 0.027 0.712 0.417 0.25 0.44 0.405 0.301 Dataset Mixture Model Dataset Mixture MLQA $(\uparrow)$ XLSUM $(\uparrow)$ XQuAD $(\uparrow)$ GMMLU $(\uparrow)$ Average Mistral-7B Alpaca MultiAlpaca 0.406 0.079 0.217 0.41 0.278 LLaMA-3-8B Alpaca MultiAlpaca 0.480 0.040 0.139 0.50 0.289
Table 2: Multilingual Ability results for two-phase Continual Fine-tuning (CFT). With green, we denote an improvement in Multilingual Ability post Phase 2 fine-tuning. Likewise, we denote a decline in Multilingual Ability with red. For MLQA and XQUAD we use F1 abstractive score, while for XLSUM we use ROUGE Lin (2004) score. We also provide numbers for dataset mixture – when the models are fine-tuned simultaneously on the Phase 1 and Phase 2 datasets.
We compare the English and Multilingual ability of Mistral-7B and LLaMA-3-8B continually fine-tuned models on different phase-wise datasets [^5] When it is clear from the context, we use “Instruct” to denote the dataset used in Phase 1 to instruction fine-tune Mistral-7B-Instruct or LLaMA-3-8B-Instruct.. Table 1 presents the results for English Ability (EA), while Table 2 presents the results for Multilingual Ability (MA). Table 2 reports the average score across languages. We provide language-specific scores and results when the phases are reversed (e.g., MultiAlpaca-Alpaca) in §B and §C.
Comparison with Mixture.
From Tables 1 & 2: for Mixture, the mean of EA and MA scores for Mistral-7B fine-tuned on Alpaca-MultiAlpaca is 0.325, and 0.312 for LLaMA-3-8B. The corresponding two-phase mean score is 0.355 for Mistral-7B and 0.346 for LLaMA-3-8B. Our two-phase CFT setup is more effective than Mixture, for approximately the same number of training steps.
Discussion.
From Table 1, for phase-wise datasets like Instruct and MultiAlpaca, the performance of the Phase 2 models trained on them declines for English. This decline occurs when they are continually fine-tuned on multilingual data in Phase 2. However, we see a jump in Mistral-7B’s multilingual ability for the multilingual generative tasks (Table 2). That is, Phase 2 models fine-tuned on multilingual datasets show forgetting in English. However, for phase-wise datasets like Alpaca followed by MultiAlpaca, we see that Phase 2 models do not show a decline in English ability (Table 1). We also see a gain in these models’ multilingual ability (Table 2).
Ablations.
In Tables B2 & B3 (§B), we present results for OpenOrca-mOpenOrca phase-wise datasets. First, the "dataset mixture" again performs worse on average than CFT: 0.186 vs. 0.372 for Mistral-7B and 0.217 vs. 0.366 for LLaMA-3-8B. Second, for Mistral-7B, the average English ability of the Phase 2 model (over Phase 1’s Mistral-7B-OpenOrca) marginally improves: 0.413 from 0.407. Whereas, for Mistral-7B-Instruct, the average decline in English ability is significant: 0.30 from 0.450. Likewise, for LLaMA-3-8B, the average English ability for LLaMA-3-8B OpenOrca mOpenOrca sees an increase to 0.356 from 0.337. In contrast, for Instruct-mOpenOrca, the English ability significantly drops, from 0.372 to 0.138.
Observation.
With Table 1, we see that our two-phase CFT setup for multilingual adaptation shows an interesting trend: for certain pairs of phase-wise datasets (e.g., Alpaca & MultiAlpaca), the LLM after Phase 2 sees an improvement in the English ability (computed on English evaluation tasks). We notice that phase-wise datasets like Alpaca and MultiAlpaca have the same seed prompts. Alternately, the two datasets encode the same instructions in different languages. We hypothesize an LLM fine-tuned on either of these datasets learns the same instructions, and therefore, the second phase of CFT leads to lesser interference in the representation space. That is, an LLM continually fine-tuned on Alpaca & MultiAlpaca preserves its English ability across phases. We next define two metrics that aim to quantify the instruction-specific similarity of two datasets.
4.3 Similarity of Phase-wise Datasets
Phase 1 Dataset Phase 2 Dataset DES ($\uparrow$) Alpaca MultiAlpaca 0.924 mOpenOrca 0.792 OpenOrca mOpenOrca 0.953 MultiAlpaca 0.774 Mistral-7B Instruct MultiAlpaca 0.746 : Prepared using model responses on MTBench Zheng et al. (2024)
Table 3: Quantifying Phase-wise Dataset Similarity using DES: higher the score, greater the dataset similarity.
Dataset Embedding Similarity (DES).
To quantify whether two datasets are similar [^6] The CL-ML literature often defines task similarity via permutation tasks, emphasizing input-output transformations Goldfarb et al. (2024). Whereas, we consider semantic and structural similarity in natural language instructions., we define DES that computes a similarity score using the dot product of the average representations (embeddings) generated by a language-agnostic model.
Definition 1 (Dataset Embedding Similarity (DES)).
Given a language-agnostic text embedding model $\Theta$, and any pair of datasets $D_{1}$ and $D_{2}$, let DES be the function $f_{\textsf{DES}}:D\times D\rightarrow[0,1]$
$f_{\textsf{DES}}(D_{1},D_{2};\Theta)=\langle\mathbf{E}_{\Theta}(D_{1}),\mathbf{E}_{\Theta}(D_{2})\rangle$
Here, $\mathbf{E}_{\Theta}(D_{i})\in\mathbb{R}^{d},~\forall i\in\{1,2\}$ is the normalized mean embedding across samples in $D_{i}$.
Higher the DES score, more similar the embedding, i.e., greater similarity between $D_{1}$ and $D_{2}$. For $\Theta$, we use the language-agnostic sentence-tokenizer LaBSE Feng et al. (2020). We compute DES by encoding 500 random samples from Alpaca, MultiAlpaca, OpenOrca, and mOpenOrca, and measure $f_{\textsf{DES}}$ for each pair. Table 3 presents the numbers. For dataset pairs with similar datasets, we see a high DES score and relatively low scores for dissimilar datasets. DES captures the (pair-wise) variation in instruction similarity of these datasets.
Dataset $D_{2}$ Model Parameter Difference ($\downarrow$) Alpaca 0.29 Instruct 1.00 OpenOrca 0.55
Table 4: Quantifying Phase-wise Similarity using MPD: lower the score, greater the dataset similarity. Here, we fix MultiAlpaca as $D_{1}$ and $\theta_{B}$ as Mistral-7B.
Model Parameter Difference (MPD).
Another method of quantifying the similarity of instructions for two datasets $D_{1}$ and $D_{2}$ is to compute the difference between the parameters of models $\Theta_{1}$ (fine-tuned on $D_{1}$) and $\Theta_{2}$ (fine-tuned on $D_{2}$). Geometrically, the difference of the parameters captures the representation shift of $\Theta_{2}$ in the space defined by $\Theta_{1}$. If $D_{1}$ & $D_{2}$ encode the same datasets, the combined shift by $\Theta_{2}$ should be relatively lower, compared to the shift if $D_{1}$ & $D_{2}$ encode different intstructions. Formally,
Definition 2 (Model Parameter Difference (MPD)).
Given any two models $\Theta_{1}$ and $\Theta_{2}$ fine-tuned on self-instruct datasets $D_{1}$ and $D_{2}$ respectively, from the same base model $\Theta_{B}$, let MPD be the function $f_{\textsf{MPD}}:\Theta\times\Theta\rightarrow\mathbb{R}_{\geq 0}$ s.t.
$f_{\textsf{MPD}}(\Theta_{1},\Theta_{2};\Theta_{B})=\frac{1}{n}\sum_{i=1}^{n}\|\mathbf{w}(\Theta_{1,i})-\mathbf{w}(\Theta_{2,i})\|_{2}$
Here, $\mathbf{w}(\Theta_{j,i}),~\forall j\in\{1,2\}$ is $\Theta_{j}$’s $i\textsuperscript{th}$ parameter.
The smaller the MPD score, the closer the fine-tuned models are in the parameter space. Fixing Mistral-7B as the base model $\Theta_{B}$, and $D_{1}$ as MultiAlpaca, we vary $D_{2}$ as one of AlpacaOpenOrca, and mOpenOrca, and observe the corresponding MPD scores. We normalize the MPD scores with the maximum observed score across all three models for a fair comparison (see Table 4). MPD shows a similar trend to DES: for Alpaca and MultiAlpaca, the scores are lower, highlighting the similarities in the datasets in the parameter space. We see relatively higher scores for the other pair of models, implying a difference in the dataset pairs.
4.4 Visualizing Decline in English Ability
Setup.
To explain the effect of similar phase-wise data sets on an LLM’s EA, we look at model representations when parsing English. We feed MTBench Zheng et al. (2024) to the models, a widely-used English benchmark for generalized instruction-following evaluation, and visualize the similarity between the mean hidden activations for each model layer. For the analysis, given an LLM $\Theta$ with $l$ layers, let $X_{\Theta}\in\mathbb{R}^{l\times d}$ be the mean hidden activations, across $n$ samples from MTBench.
t-SNE Visualization.
Figure 1 depicts t-SNEs van der Maaten and Hinton (2008) for $X_{\textsc{Mistral-7B}}$ and $X_{\textsc{LLaMA-3-8B}}$ when these are continually fine-tuned on (i) Alpaca & MultiAlpaca and (ii) Instruct & MultiAlpaca. We observe that for similar phase-wise datasets, the model before and after Phase 2 produces similar hidden activations. Contrarily, for non-similar phase-wise datasets, the hidden activations form distinct clusters, implying separation between the phase-wise activations. That is, the model representations for non-similar phase-wise datasets are well-separated. The separation between model representations results in increased weight interference during Phase 2 – leading to a decline in EA.
Visualizing Variance in Model Representations.
Figure 1 provides an intuition for the correlation between phase-wise datasets and the decline in English ability. To further understand the layer-wise behavior of the hidden activations, similar to Chang et al. (2022), we compute covariance matrices $\Sigma_{\Theta}$ for each $X_{\Theta}$. Intuitively, $\Sigma_{\Theta}$ captures the variance in different directions for representations of hidden activations for $\Theta$.
Figure 2: Visualizing variance in model representations. We plot $\|\Sigma_{\text{Phase 2}}-\Sigma_{\text{Phase 1}}\|_{2}$ across layers to measure representation shift. Dissimilar datasets (e.g., Instruct & MultiAlpaca) induce substantially larger variation than similar ones (e.g., Alpaca& MultiAlpaca). For LLaMA-3-8B, this variation is consistently high across layers, whereas for Mistral-7B, it is more localized and increases primarily in higher layers.
We first compute the mean centered activation matrix $\bar{X}_{\Theta}=X_{\Theta}-\mu_{\Theta}$, where $\mu_{\Theta}=1/l\sum_{i=1}^{l}X_{\Theta}^{(i)}$. Next, we derive $\Sigma_{\Theta}=\frac{1}{l-1}\cdot\bar{X}_{\Theta}^{T}\bar{X}_{\Theta}\in\mathbb{R}^{d\times d}$. To compare the layer-wise variance in representations, we compute the L2-Norm of the difference of the matrices $\Sigma_{\textsc{Mistral-7B}}$ (Figure 2 (left)) or $\Sigma_{\textsc{LLaMA-3-8B}}$ (Figure 2 (right)) when continually fine-tuned on Alpaca & MultiAlpaca (blue lines) or Instruct & MultiAlpaca (red lines).
From the figures, we see clear evidence of representational change, both in terms of the magnitude of the change and the subset of layers that show a greater change. For Mistral-7B, the Phase 2 model after CFT with Instruct & MultiAlpaca, shows 3 to 4 times more variation in its representations compared to the model with Alpaca & MultiAlpaca phase-wise datasets. This gap is significantly larger for LLaMA-3-8B.
CFT Setup English Ability (EA) Multilingual Ability (MA) Combined Mitigating IFEval Alpaca Eval MMLU HellaSwag XLSUM_en Avg MLQA XLSUM XQUAD GMMLU Avg Avg Strategy $(\uparrow)$ $(\uparrow)$ $(\uparrow)$ $(\uparrow)$ $(\uparrow)$ $(\uparrow)$ $(\uparrow)$ $(\uparrow)$ $(\uparrow)$ $(\uparrow)$ $(\uparrow)$ $(\uparrow)$ Mistral-7B 0.462 0.15 0.533 0.416 0.10 0.332 0.307 0.033 0.436 0.430 0.302 0.317 LF_H1 0.456 0.03 0.497 0.598 0.10 0.336 0.176 0.016 0.215 0.428 0.209 0.273 LF_H2 0.364 0.12 0.364 0.504 0.12 0.294 0.213 0.014 0.442 0.384 0.263 0.279 Spectrum 0.435 0.24 0.488 0.524 0.13 0.363 0.317 0.083 0.176 0.370 0.237 0.30 GR_5 0.540 0.17 0.540 0.611 0.11 0.394 0.311 0.008 0.428 0.445 0.298 0.346 GR_10 0.567 0.12 0.567 0.594 0.12 0.394 0.213 0.007 0.427 0.450 0.274 0.334 ER_10 0.593 0.08 0.580 0.635 0.13 0.404 0.249 0.008 0.398 0.448 0.276 0.340 LoRA 0.383 0.09 0.579 0.625 0.03 0.341 0.289 0.043 0.518 0.435 0.321 0.331 LLaMA-3-8B 0.182 0.10 0.239 0.278 0.09 0.178 0.321 0.030 0.417 0.440 0.302 0.240 LF_H1 0.303 0.0 0.231 0.275 0.01 0.164 0.368 0.037 0.505 0.237 0.287 0.225 LF_H2 0.380 0.06 0.485 0.525 0.08 0.306 0.400 0.038 0.505 0.338 0.320 0.313 Spectrum 0.409 0.09 0.612 0.524 0.01 0.329 0.429 0.056 0.086 0.472 0.261 0.295 GR_5 0.269 0.01 0.516 0.316 0.07 0.236 0.437 0.019 0.593 0.342 0.348 0.292 GR_10 0.264 0.12 0.229 0.250 0.0 0.173 0.254 0.009 0.314 0.238 0.204 0.189 ER_10 0.420 0.02 0.603 0.561 0.12 0.345 0.434 0.025 0.53 0.448 0.359 0.352 LoRA 0.196 0.0 0.280 0.235 0.0 0.142 0.007 0.008 0.005 0.278 0.075 0.109
Table 5: English Ability (EA) and Multilingual Ability (MA) results for our mitigating strategies. These comprise Generative Replay (GR_5 & GR_10), English Replay (ER_10) and Layer Freezing (LF_H1, LF_H2 & Spectrum). We use LoRA Hu et al. (2022) as a baseline strategy. For ER_10, we use the English dataset used in GR with original responses. The Phase 1 dataset is Instruct for each row, while Phase 2 is MultiAlpaca. The first row for both Mistral-7B and LLaMA-3-8B provides numbers for Instruct-MultiAlpaca (from Table 1 &  2).
5 Mitigating Strategies for CFT
To mitigate EA decline, we explore two tailored CFT techniques [^7] In §F, we discuss how MAD-X Pfeiffer et al. (2020) and other PEFT-based methods Badola et al. (2023) are not suitable for our setting.: Distribution Replay and Layer Freezing. In Distribution Replay, we study Generative Replay (GR), a new English data generation method inspired by dataset similarity and English ability (§4.2), and English Replay (ER), which replays parallel English data of Phase 2’s distribution. In Layer Freezing (LF), we identify layers to freeze during Phase 2 fine-tuning using specific heuristics.
5.1 Distribution Replay
Typically, Generative Replay (GR) is a technique that generates data from past distributions to be used alongside new task data for the continual fine-tuning of a model on a new task Shin et al. (2017). However, from §4.2, we do not see a decline in English ability if the phase-wise datasets encode similar instructions. Based on this, we use the Phase 1 model to generate responses, in English, from the English counterpart of the multilingual dataset used for fine-tuning in Phase 2. The intuition is that the generated dataset may bridge the distributions of Phase 1 and Phase 2.
During Phase 2 fine-tuning, we include varying quantities of this generated data: specifically, 5% (GR_5) and 10% (GR_10), of the Phase 2 dataset. We also fine-tune the models with a similar sized subset of the English counterpart with original responses [^8] This dataset may not be available for all multilingual datasets, such as Aya Singh et al. (2024b). While instructions can be translated into English, translating responses is often impractical. Thus, ER is the best-case scenario for GR.. We refer to this mitigating strategy as English Replay (ER_10).
5.2 Layer Freezing
Model regularization is an effective technique to mitigate the drop in the previous task’s performance in continual learning (e.g., EWC Kirkpatrick et al. (2017)). However, this is computationally inefficient as it requires using both the old and new sets of parameters. Instead, we use Layer Freezing (LF), a relatively efficient technique for use as a ‘regularizer’ to preserve English ability during Phase 2. We consider the following variations to select the set of layers to freeze. These variants allow us to study how different criteria for selecting frozen layers affect the trade-off between preserving English ability and enabling multilingual adaptation.
  1. 1.
    LF_H1: freezing a random set of 10 layers of the model from Phase 1 to be fine-tuned in Phase 2.
  2. 2.
    LF_H2: freezing the top-10 layers that have changed the most during Phase 1 fine-tuning (e.g., Mistral-7B Base to Mistral-7B-Instruct). We select layers separately for Key, Query, and Value, for each attention head.
  3. 3.
    Spectrum Hartford et al. (2024): freeze the "most informative" layers of the Phase 1 model based on their signal-to-noise ratio (refer to §D.1).
We present our results in Table 5 for both GR and LF. We define a baseline in which we use LoRA Hu et al. (2022) [^9] Parameter efficient techniques like LoRA Hu et al. (2022) are also widely used to efficiently fine-tune LLMs on multilingual data. However, such techniques also show forgetting on English Aggarwal et al. (2024) after Phase 2.for continually fine-tuning in Phase 2. We perform LoRA fine-tuning with rank 64 and quantisation bfloat16.
5.3 Results Discussion
From Table 5, we see that GR, ER, and LF mitigate the decline in EA and also show gains in MA.
Distribution Replay.
ER_10 demonstrates the best performance in both English and combined ability, with EA scores of 0.404 for Mistral-7B and 0.345 for LLaMA-3-8B, and the best combined average. GR_5 also excels in multilingual tasks, performing similar to ER_10: 0.298 vs. 0.276 for Mistral-7B and 0.348 vs. 0.359 for LLaMA-3-8B. GR_5 also performs reasonably well on English tasks, achieving scores of 0.394 and 0.236 for Mistral-7B and LLaMA-3-8B, respectively, making it a competitive strategy.
Layer Freezing.
Compared to ER and GR, LF_H1, LF_H2, and Spectrum show mixed results. LF_H2 performs better than LF_H1. Spectrum’s EA scores are better than LF_H1 and LF_H2, but suffers from lower multilingual numbers.
Additional Discussion & Results.
In §D.3, we also present EA and MA results for Mistral-7B Instruct-mOpenOrca for our mitigating strategies. Here, LF, particularly Spectrum, performs better than the other strategies. Furthermore, in §D.4, we analyze the computational cost of these strategies over the baseline CFT setup.
5.4 Controlling Representation Drift under Mitigation
Figure 3: Layer-wise covariance shift under mitigation strategies in Phase 2 CFT. We plot $\|\Sigma_{\text{Phase 2}}-\Sigma_{\text{Phase 1}}\|_{2}$ across layers. The vanilla Phase 2 model (Figure 2) exhibits a sharp increase in drift in higher layers. Replay (GR) substantially reduces the magnitude of this drift across all layers, while freezing (LF_H2) confines it to a smaller subset of layers.
Building on Figure 2, we examine how mitigation strategies affect representation drift relative to the vanilla Phase 2 baseline (Mistral-7B Instruct $\rightarrow$ MultiAlpaca). As in §4.4, we measure drift as $\|\Sigma_{\text{Phase 2}}-\Sigma_{\text{Phase 1}}\|_{2}$, where larger values indicate greater deviation from Phase 1 representations.
Consistent with earlier observations, the mitigation strategies explicitly designed to curb representational change (LF & GR) exhibit lower drift than the baseline model across layers. Replay-based methods (GR) reduce the overall magnitude of drift while preserving its layer-wise structure, whereas freezing-based methods (LF_H2) constrain the drift to a subset of layers.
Notably, generative replay remains closest to the Phase 1 representation geometry among all strategies. This closer alignment correlates with improved task and language performance compared to the vanilla Phase 2 model (Table 1 and Table 2). These results reinforce our central hypothesis: in two-phase CFT with dissimilar datasets, representation shift is the primary driver of English degradation, and explicitly controlling this shift preserves alignment with Phase 1 representations, leading to improved English retention without sacrificing multilingual gains.
6 Conclusion & Future Work
We study the role of dataset similarity in a novel two-phase continual fine-tuning (CFT) for multilingual adaptation. Across Mistral-7B and LLaMA-3-8B, we show that alignment between Phase 1 and Phase 2 datasets is critical: similar datasets preserve English ability while improving multilingual performance, whereas misaligned datasets lead to degradation. We identify representation shift as the underlying mechanism driving this behavior. In particular, misaligned fine-tuning induces large covariance shifts in hidden representations, concentrated in higher layers. We show that mitigation strategies such as generative replay and layer freezing improve English retention by controlling this drift, either by globally reducing its magnitude or by constraining it to specific layers, while preserving multilingual gains.
Future Work.
Our results suggest several directions for future research. First, designing adaptive mitigation strategies that dynamically control representation drift during training could provide stronger and more efficient trade-offs than fixed heuristics Zhang et al. (2026). Second, developing principled measures of instruction-level dataset similarity may enable better prediction and prevention of degradation. Finally, exploring parameter-efficient approaches that explicitly regularize representation geometry could reduce the computational overhead of current methods while maintaining their effectiveness.
7 Limitations
Our study has several limitations. First, we rely on DES and MPD as proxies for dataset similarity; while effective in our setting, these metrics may not capture all nuances of instruction-level similarity. Second, our experiments are limited to Mistral-7B and LLaMA-3-8B, and the observed trends may not fully generalize to models with different architectures, scales, or training paradigms. Similarly, the effectiveness of our mitigation strategies (i.e., replay and layer freezing) may vary across datasets and model families. Third, our experimental setup is constrained by computational resources. We were unable to explore larger models, longer training schedules, or broader hyperparameter sweeps, which may further influence the observed trade-offs. Fourth, some mitigation strategies impose additional requirements. In particular, the best-performing method, ER_10, assumes access to parallel data, which may not always be available in practice. Finally, our evaluation focuses on specific benchmarks for task and language ability. While these provide controlled comparisons, they may not fully capture real-world performance across diverse applications.
Appendix A Training Details
A.1 Hyperparameters for Fine-tuning and Training Setup
Hyperparameter Value Learning Rate $1\times 10^{-6}$ Epochs 4 Global Batch size 16 Scheduler Cosine Warmup Linear Warmup Steps 10 Optimizer AdamW Loshchilov and Hutter (2019) Weight Decay 0
Table A1: Hyperparameters for our Two-phase Continual Fine-tuning (CFT)
A.2 Fine-tuning Datasets
OpenOrca is an English-only self instruct dataset, created to best mimic the Orca dataset Mukherjee et al. (2023), which is not publicly available. To create the multilingual version of OpenOrca, namely mOpenOrca, we follow Ahuja et al. (2024b) to generate selective translations for a subset of OpenOrca. The subset contains 50k samples from the OpenOrca dataset and we selectively translate them to 11 languages which are also in MultiAlpaca. In total, we generate 550k examples for all languages.
A.3 Evaluation Tasks
In this paper, we consider two sets of benchmarks to evaluate task and language ability. We explain them briefly next.
English Ability (EA).
To quantify an LLM’s task ability, we evaluate Phase 1 and Phase 2 models on the following tasks:
  1. 1.
    IFEval Zhou et al. (2023): Instruction-Following Evaluation (IFEval) asses the ability of an LLM to follow natural language instructions. It comprises 500 verifiable instructions (e.g., “mention the keyword AI 3 times”). We choose IFEval as the instructions are verifiable and also test an LLM’s context understanding.
  2. 2.
    Alpaca Eval Li et al. (2023): This is an LLM-based automatic evaluator for instruction following models, to measure task ability. Like Aggarwal et al. (2024), we evaluate our CFT models against text-davinci-003 responses on 800 instructions and use GPT4 (gpt-4-32k) as the evaluator.
  3. 3.
    MMLU Hendrycks et al. (2021): Massive Multitask Language Understanding (MMLU) is a benchmark to assess an LLM’s knowledge and problem-solving abilities. It includes 57 subjects across domains like STEM, or law, with 16k MCQs in total.
  4. 4.
    HellaSwag Zellers et al. (2019): This is a popular benchmark to evaluate the commonsense reasoning ability of an LLM. HellaSwag’s test split contains 10k samples in total.
Multilingual Ability (MA).
To quantify an LLM’s language ability, we evaluate our fine-tuned models on three benchmark datasets comprising two multilingual generative tasks: question answering and summarization.
  • Question Answering: MLQA Lewis et al. (2019) contains 5k extractive question-answering instances in 7 languages. The XQuAD dataset Artetxe et al. (2019) consists of a subset of 240 paragraphs and 1190 question-answer pairs across $1$ languages.
  • Summarisation: XLSUM Hasan et al. (2021) spans 45 languages, and we evaluate our models in Arabic, Chinese-Simplified, English, French, Hindi, Japanese, and Spanish.
Appendix B Evaluating Multilingual Ability for Continual Fine-tuning
Two-phase Continual Fine-tuning Model Phase 1 (P1) Phase 2 (P2) IFEval $(\uparrow)$ Alpaca Eval $(\uparrow)$ MMLU $(\uparrow)$ HellaSwag $(\uparrow)$ XLSUM_en $(\uparrow)$ Average Dataset Dataset P1 P2 P1 P2 P1 P2 P1 P2 P1 P2 P1 P2 Mistral-7B OpenOrca mOpenOrca 0.494 0.482 0.31 0.32 0.601 0.582 0.612 0.562 0.02 0.12 0.407 0.413 Instruct mOpenOrca 0.550 0.426 0.35 0.06 0.575 0.507 0.641 0.509 0.13 0.0 0.450 0.30 LLaMA-3-8B OpenOrca mOpenOrca 0.377 0.425 0.09 0.07 0.579 0.599 0.571 0.564 0.07 0.12 0.337 0.356 Instruct mOpenOrca 0.735 0.205 0.14 0.0 0.340 0.236 0.533 0.250 0.11 0.0 0.372 0.138 Dataset Mixture Model Dataset Mixture IFEval $(\uparrow)$ Alpaca Eval $(\uparrow)$ MMLU $(\uparrow)$ HellaSwag $(\uparrow)$ XLSUM_en $(\uparrow)$ Average Mistral-7B OpenOrca mOpenOrca 0.228 0.035 0.284 0.444 0.02 0.202 LLaMA-3-8B OpenOrca mOpenOrca 0.248 0.072 0.484 0.473 0.0 0.255
Table B2: English Ability results for two-phase Continual Fine-tuning (CFT). With green, we highlight an increase in a model’s task ability post P2 fine-tuning. Likewise, red highlights a decline in a model’s task ability.
Two-phase Continual Fine-tuning Model Phase 1 (P1) Phase 2 (P2) MLQA $(\uparrow)$ XLSUM $(\uparrow)$ XQuAD $(\uparrow)$ GMMLU $(\uparrow)$ Average Dataset Dataset P1 P2 P1 P2 P1 P2 P1 P2 P1 P2 Mistral-7B OpenOrca mOpenOrca 0.435 0.360 0.007 0.008 0.556 0.643 0.433 0.312 0.358 0.331 Instruct mOpenOrca 0.246 0.155 0.012 0.040 0.351 0.323 0.440 0.279 0.262 0.20 LLaMA-3-8B OpenOrca mOpenOrca 0.401 0.453 0.017 0.006 0.499 0.531 0.242 0.513 0.290 0.376 Instruct mOpenOrca 0.609 0.604 0.048 0.048 0.712 0.713 0.250 0.233 0.405 0.40 Dataset Mixture Model Dataset Mixture MLQA $(\uparrow)$ XLSUM $(\uparrow)$ XQuAD $(\uparrow)$ GMMLU $(\uparrow)$ Average Mistral-7B OpenOrca mOpenOrca 0.201 0.128 0.071 0.277 0.169 LLaMA-3-8B OpenOrca mOpenOrca 0.224 0.034 0.091 0.364 0.178
Table B3: Multilingual Ability results for two-phase Continual Fine-tuning (CFT). With green, we highlight an increase in a model’s Multilingual ability post Phase 2 fine-tuning. Likewise, red highlights a decline in a model’s Multilingual ability.
English Ability.
Table B2 present the english ability numbers of our ablations on the OpenOrca-mOpenOrcaand Instruct-mOpenOrcadatasets using Mistral-7B and LLaMA-3-8B models. When the datasets are pairwise not similar, i.e., Instruct-mOpenOrca, Mistral-7B shows a significant decline in the average English ability, from 0.450 in Phase 1 to 0.30 in Phase 2. Likewise, LLaMA-3-8B also experiences a decrease, dropping from 0.372 to 0.138 on average.
In contrast, when the pairwise datasets are similar, i.e., OpenOrca and mOpenOrca, Mistral-7B sees a marginal increase between the phases $(0.407\rightarrow 0.413)$, on average. LLaMA-3-8B’s performance sees an improvement in the average English ability, from 0.337 to 0.356.
Multilingual Ability.
Table B3 tabulates the results for multilingual ability. We see an improvement in the average multilingual ability for the OpenOrca-mOpenOrca dataset pair for LLaMA-3-8B. For Mistral-7B, there is a marginal drop $(0.358\to 0.331)$. For Instruct-mOpenOrca, with LLaMA-3-8B, the average multilingual ability is virtually the same across tasks (0.405 vs. 0.40). However, for Mistral-7B, we see a slight drop in the average language ability, driven primarily due to a decline in performance for MLQA.
Model Phase 1 (P1) Phase 2 (P2) IFEval $(\uparrow)$ Alpaca Eval $(\uparrow)$ MMLU $(\uparrow)$ HellaSwag $(\uparrow)$ Average Dataset Dataset P1 P2 P1 P2 P1 P2 P1 P2 P1 P2 Mistral-7B MultiAlpaca Alpaca 0.245 0.290 0.120 0.114 0.528 0.430 0.476 0.510 0.342 0.336 LLaMA-3-8B MultiAlpaca Alpaca 0.245 0.340 0.038 0.065 0.570 0.540 0.577 0.590 0.357 0.384 Mistral-7B mOpenOrca OpenOrca 0.190 0.310 0.091 0.055 0.410 0.490 0.520 0.510 0.303 0.341 LLaMA-3-8B mOpenOrca OpenOrca 0.314 0.340 0.0 0.0 0.530 0.540 0.522 0.590 0.342 0.368
Table B4: English Ability results for two-phase Continual Fine-tuning (CFT)
Model Phase 1 Phase 2 MLQA $(\uparrow)$ XLSUM $(\uparrow)$ XQuAD $(\uparrow)$ Average Dataset Dataset Phase 1 Phase 2 Phase 1 Phase 2 Phase 1 Phase 2 Phase 1 Phase 2 Mistral-7B MultiAlpaca Alpaca 0.122 0.230 0.021 0.030 0.122 0.090 0.088 0.116 LLaMA-3-8B MultiAlpaca Alpaca 0.363 0.340 0.048 0.040 0.058 0.030 0.157 0.134 Mistral-7B mOpenOrca OpenOrca 0.165 0.160 0.077 0.070 0.140 0.180 0.127 0.137 LLaMA-3-8B mOpenOrca OpenOrca 0.057 0.0 0.038 0.0 0.047 0.0 0.047 0.0
Table B5: Multilingual Ability results for two-phase Continual Fine-tuning (CFT)
Furthermore, Table B6, Table B7, and Table B8 present the language-specific results for MLQA, XLSUM, and XQuAD, respectively.
Model Phase 1 Phase 2 MLQA Dataset Dataset Phase 1 Phase 2 ar de es hi vi zh ar de es hi vi zh Mistral-7B Alpaca MultiAlpaca 0.143 0.337 0.331 0.149 0.385 0.031 0.172 0.485 0.529 0.196 0.336 0.009 Instruct 0.113 0.440 0.395 0.088 0.369 0.073 0.228 0.456 0.529 0.279 0.327 0.0222 LLaMA-3-8B Alpaca 0.320 0.538 0.563 0.438 0.611 0.155 0.552 0.672 0.765 0.573 0.784 0.237 Instruct 0.549 0.701 0.769 0.624 0.788 0.192 0.316 0.453 0.526 0.137 0.464 0.028 Mistral-7B OpenOrca mOpenOrca 0.374 0.504 0.511 0.395 0.600 0.226 0.298 0.506 0.572 0.274 0.481 0.030 Instruct 0.113 0.440 0.395 0.088 0.369 0.073 0.115 0.253 0.213 0.088 0.222 0.038 LLaMA-3-8B OpenOrca 0.262 0.545 0.565 0.369 0.568 0.099 0.437 0.549 0.622 0.462 0.625 0.024 Instruct 0.320 0.538 0.563 0.438 0.611 0.155 0.554 0.701 0.771 0.625 0.787 0.188
Table B6: MLQA: Language Ability results for two-phase Continual Fine-tuning (CFT).
Model Phase 1 Phase 2 XLSUM Dataset Dataset Phase 1 Phase 2 Arabic Chinese_simplified french Hindi Japanese Spanish Arabic Chinese_simplified french Hindi Japanese Spanish Mistral-7B Alpaca MultiAlpaca 0.001 0.012 0.025 0.001 0.012 0.023 0.022 0.034 0.112 0.016 0.067 0.106 Instruct 0.001 0.005 0.028 0.001 0.009 0.025 0.016 0.015 0.060 0.010 0.040 0.056 LLaMA-3-8B Alpaca 0.005 0.015 0.071 0.003 0.037 0.067 0.003 0.018 0.073 0.002 0.041 0.070 Instruct 0.008 0.015 0.092 0.004 0.080 0.087 0.002 0.013 0.055 0.001 0.055 0.051 Mistral-7B OpenOrca mOpenOrca 0.001 0.010 0.014 0.001 0.007 0.009 0.001 0.006 0.018 0.001 0.008 0.016 Instruct 0.001 0.005 0.028 0.001 0.009 0.025 0.007 0.017 0.092 0.005 0.030 0.088 LLaMA-3-8B OpenOrca 0.000 0.003 0.061 0.000 0.004 0.035 0.000 0.003 0.016 0.001 0.000 0.013 Instruct 0.008 0.015 0.092 0.004 0.080 0.087 0.007 0.015 0.091 0.004 0.082 0.087
Table B7: XLSUM: Language Ability results for two-phase Continual Fine-tuning (CFT).
Model Phase 1 Phase 2 XQuAD Dataset Dataset Phase 1 Phase 2 ar de el es hi ro ru th tr vi zh ar de el es hi ro ru th tr vi zh Mistral-7B Alpaca MultiAlpaca 0.194 0.379 0.248 0.374 0.224 0.418 0.150 0.185 0.454 0.475 0.088 0.613 0.692 0.657 0.713 0.670 0.679 0.661 0.385 0.666 0.734 0.148 Instruct 0.166 0.568 0.260 0.510 0.173 0.508 0.336 0.210 0.460 0.502 0.168 0.369 0.612 0.253 0.634 0.450 0.553 0.555 0.180 0.532 0.566 0.089 LLaMA-3-8B Alpaca 0.393 0.689 0.529 0.735 0.644 0.723 0.538 0.398 0.671 0.748 0.376 0.676 0.850 0.710 0.893 0.740 0.817 0.726 0.526 0.770 0.884 0.519 Instruct 0.659 0.795 0.702 0.852 0.715 0.810 0.609 0.594 0.728 0.834 0.533 0.444 0.580 0.244 0.657 0.241 0.586 0.493 0.092 0.580 0.558 0.113 Mistral-7B OpenOrca mOpenOrca 0.001 0.010 0.014 0.001 0.007 0.009 0.001 0.006 0.018 0.001 0.008 0.639 0.832 0.570 0.847 0.601 0.776 0.771 0.366 0.734 0.820 0.113 Instruct 0.166 0.568 0.260 0.510 0.173 0.508 0.336 0.210 0.460 0.502 0.168 0.256 0.457 0.320 0.443 0.256 0.409 0.215 0.245 0.364 0.428 0.162 LLaMA-3-8B OpenOrca 0.505 0.642 0.587 0.711 0.604 0.634 0.651 0.290 0.699 0.685 0.104 0.639 0.832 0.570 0.847 0.601 0.776 0.771 0.366 0.734 0.820 0.113 Instruct 0.659 0.795 0.702 0.852 0.715 0.810 0.609 0.594 0.728 0.834 0.533 0.654 0.793 0.703 0.852 0.718 0.808 0.606 0.600 0.729 0.836 0.540
Table B8: XQuAD: Language Ability results for two-phase Continual Fine-tuning (CFT).
Appendix C Reverse Order CFT Result Analysis
In Tables B4 and B5, we reverse the order of Phase 1 and Phase 2 datasets, first fine-tuning on the multilingual dataset followed by its English counterpart. For Mistral-7B (MultiAlpaca-Alpaca), the average performance is 0.226, and for LLaMA-3-8B (MultiAlpaca-Alpaca), it is 0.259. Compared to the mixture setting and the Alpaca-MultiAlpaca configuration (§4), we observe that English ability benefits from multilingual fine-tuning in Phase 1, resulting in performance comparable to the data mixture setting. However, fine-tuning on English data in Phase 2 leads to a drastic drop in multilingual ability, yielding worse results than both the mixture setting and the two-phase setup considered in the main paper.
Appendix D Mitigating Strategies
Here, we provide additional details on Spectrum Hartford et al. (2024). We then visualize the impact of our mitigating strategies on the variance in model representations. Lastly, we ablate our findings for the Instruct-mOpenOrca phase-wise datasets.
D.1 Spectrum
Spectrum Hartford et al. (2024) is a layer-freezing technique that optimizes the fine-tuning of LLMs by selecting layers based on their signal-to-noise ratio (SNR). We use Spectrum as a heuristic for layer-freezing; that is, the layers identified as "important" by Spectrum are frozen during Phase 2 fine-tuning. A layer is important based on its signal-to-noise (SNR) ratio. In the following, we elaborate on how Spectrum computes SNR.
Marchenko-Pastur distribution.
The Marchenko-Pastur distribution Marchenko and Pastur (1967) is given by:
$\rho(\lambda)=\frac{1}{2\pi\sigma^{2}}\frac{\sqrt{(\lambda_{+}-\lambda)(\lambda-\lambda_{-})}}{\lambda},$
where
$\lambda_{\pm}=\sigma^{2}(1\pm\sqrt{Q})^{2},$
and $Q=\frac{N}{M}$, with $N$ and $M$ being the dimensions of a random matrix $W$, and $\sigma^{2}$ representing the variance of the entries in $W$.
SNR.
Let $W\in\mathbb{R}^{N\times M}$ be the weight matrix of a given layer. The empirical spectral density of $W$ is analyzed by comparing its eigenvalue distribution of $1/N\cdot W^{T}W$ against the theoretical Marchenko-Pastur distribution. Deviations from this distribution indicate the presence of significant signal components. We get,
$\lambda_{\pm}=\sigma^{2}\left(1\pm\sqrt{\frac{M}{N}}\right)^{2},$
where $\lambda_{\pm}$ are the largest and smallest eigenvalues and $\sigma$ the standard deviation. This implies the bounds of singular values of $W$ as:
$\epsilon_{\pm}=\frac{1}{\sqrt{N}}\sigma\left(1\pm\sqrt{\frac{M}{N}}\right)$ (1)
By evaluating how the singular values of $W$ distribute relative to $\epsilon_{\pm}$, Spectrum assesses the SNR of each layer, as defined next.
Ratio
Hartford et al. (2024). Specifically, the SNR value of a weight matrix is,
$\text{SNR}=\frac{\sum_{k|\sigma_{k}>\epsilon}\sigma_{k}}{\sum_{k|\sigma_{n}<\epsilon}\sigma_{n}}$
Here, $\epsilon$ separates signal from noisy singular values. Layers with singular values significantly exceeding $\epsilon_{+}$ have a high SNR, indicating a substantial presence of informative signal components.
Measuring the Ratio
Hartford et al. (2024). Having defined all ingredients above, Spectrum now computes each layer’s SNRs. To do this, it first computes SVD Zhang and Xu (2009) of the the layer’s weight matrix, calculates the SNR and normalizes it by the highest singular value. Eq. 1 gives the noise threshold.
Now, Spectrum selects layers with higher SNRs, where the number of layers selected is a hyperparameter. Similar to Hartford et al. (2024), for our experiments, we select the top-50% of layers in each module.
D.2 LLaMA-3-8B Doesn’t Show Consistent Improvement with our Mitigation Strategies
From Table 5, while both GR and LF improve on the baseline LLaMA-3-8B-Instruct MultiAlpaca, the gains in task and multilingual ability are not comparable to LLaMA-3-8B-Instruct.
(a) Mistral-7B
(b) LLaMA-3-8B
Figure D4: Demonstrating extent of cross-lingual transfer in Mistral-7B and LLaMA-3-8B on a parallel dataset prepared by subsampling Flores Costa-jussà et al. (2022). We find that the English activation cluster for LLaMA-3-8B is separated from the multilingual cluster, compared to Mistral-7B.
To understand this further, for GR, we investigate the cross-linguality difference between LLaMA-3-8B and Mistral-7B. Like Figure 1, we plot t-SNEs of the mean model activations for the Mistral-7B and LLaMA-3-8B base models on two parallel datasets, English and Multilingual. We create the parallel datasets by subsampling data from Flores Costa-jussà et al. (2022). In Figure D4, we see that the English activation cluster for LLaMA-3-8B is separated out from multilingual cluster, compared to Mistral-7B. This suggests that GR may not be as effective when the model has less cross lingual ability. While for LF, we acknowledge that our method to identify the layers to freeze may not be the best and better methods to identify which layers to freeze can be a direction for future work.
Last, but not the least, we acknowledge that LLaMA-3-8B-Instruct seems to be a strong model even on multilingual benchmarks. Hence, it is also important to evaluate Phase 1 models on these benchmarks first and then decide if the Phase 2 fine-tuning step should be undertaken or not.
With regards to LLaMA-3-8B-Instruct MultiAlpaca MA results in Table 2, we believe that this is due to lack of cross-linguality in LLaMA-3-8B-Instruct and less data in MultiAlpaca which fails to cause sufficient representation drift to improve the model’s performance.
D.3 Additional Ablations
We also present the impact of our mitigating strategies for the Instruct-mOpenOrca phase-wise datasets on Mistral-7B. Table D9 presents these results.
We see that LF_H2 achieves moderate success, especially in maintaining the language ability for MLQA (0.258) and XQUAD (0.527). However, task ability shows some decline (e.g., IFEval (0.401) and Alpaca Eval (0.048)), compared to the baseline. Furthermore, GR_5 results in lower task ability (IFEval = 0.281), while GR_10 performs slightly better in task ability (e.g., MMLU = 0.483, HellaSwag = 0.494). Among the baselines, ER_10 performs similarly to the generative replay strategies, with modest improvements in task ability (e.g., IFEval = 0.367, MMLU = 0.479), but still struggles in language ability. Perhaps LoRA shows the best overall performance among the strategies for maintaining task ability (e.g., IFEval = 0.587, MMLU = 0.567, HellaSwag = 0.591) with reasonable retention of language ability (e.g., XQUAD = 0.354).
These results show that no single strategy is perfect, and future work may need to combine these strategies or develop new approaches to address the balance between task and language ability retention across phases.
CFT Setup Task Ability Language Ability Overall Model Phase 2 Mitigating IFEval Alpaca Eval MMLU HellaSwag Avg MLQA XLSum XQUAD Avg Avg Dataset Strategy Mistral-7B mOpenOrca 0.426 0.060 0.507 0.509 0.376 0.155 0.040 0.323 0.173 0.275 LF_H2 0.401 0.048 0.518 0.487 0.364 0.258 0.060 0.527 0.282 0.323 Spectrum 0.442 0.158 0.508 0.616 0.431 0.387 0.086 0.201 0.225 0.328 GR_5 0.281 0.027 0.478 0.495 0.320 0.167 0.042 0.305 0.171 0.246 GR_10 0.305 0.013 0.483 0.494 0.324 0.150 0.038 0.238 0.142 0.233 ER_10 0.367 0.025 0.479 0.493 0.341 0.157 0.042 0.305 0.168 0.255 LoRA 0.587 0.130 0.567 0.591 0.469 0.167 0.027 0.354 0.183 0.326
Table D9: English and Multilingual Ability results for our mitigating strategies, Generative Replay (GR_5 & GR_10), English Replay (ER_10) and Layer Freezing (LF_H1, LF_H2 & Spectrum). We use LoRA Hu et al. (2022) as a baseline strategy. For ER_10, we use the English dataset used in GR with original responses. The Phase 1 dataset is Instruct for each row. The first row provides Mistral-7B numbers for Instruct-mOpenOrca (from Table B2).
D.4 Compute Analysis
The computational overhead of replay arises from an increase in the effective training set size . In ER_10, each epoch includes the original Phase 2 data along with an additional 10% replay buffer from Phase 1, resulting in a proportional increase in the number of training tokens and hence compute. Similarly, GR_5 augments each epoch with 5% generated replay data, leading to a smaller overhead.
In contrast, layer freezing (LF) significantly reduces computation by updating only a subset of model parameters. By freezing 50% of the layers, LF yields a corresponding reduction in training cost while still achieving a reasonable trade-off between English retention and multilingual adaptation.
Appendix E Resources Used
We conduct all experiments on 4 NVIDIA A100 GPUs (80GB each) with a 96-core AMD CPU. A single fine-tuning run on MultiAlpaca takes approximately 4 hours, while training on mOpenOrca requires around 12 hours. The models used in our experiments, along with their corresponding checkpoints and licenses, are listed below:
Appendix F Existing Mitigating Strategies
We now discuss why MAD-X Pfeiffer et al. (2020) and other PEFT-based methods Badola et al. (2023) are not well-suited to our setting.
  • Parameter-Efficient Fine-tuning for Robust Continual Multilingual Learning Badola et al. (2023): This work considers a fixed downstream task (text classification) that is incrementally extended to new languages, mitigating forgetting via task- and language-specific adapters. Their approach assumes continued access to prior task data and the ability to maintain multiple adapter checkpoints. In contrast, our setting involves end-to-end instruction fine-tuning over a heterogeneous mixture of tasks. We do not assume access to Phase 1 data during Phase 2, nor do we maintain multiple parameter sets due to the memory overhead of storing additional model weights (§3).
  • MAD-X Pfeiffer et al. (2020): MAD-X introduces bottleneck adapters within encoder-only or encoder–decoder Transformer architectures to enable cross-lingual transfer. However, recent work Zhao et al. (2024) shows that such adapters do not integrate cleanly with decoder-only LLMs (e.g., Mistral-7B, LLaMA-3-8B), which form the basis of our experiments. As our models lack an encoder stack, the routing and merging mechanisms central to MAD-X are not directly applicable.
  • Preserving Cross-Linguality via Continual Learning Liu et al. (2021): This method extends Gradient Episodic Memory by storing exemplars (or their gradients) from previous phases and replaying them during training. Our setup (§3) explicitly disallows (i) reusing Phase 1 data and (ii) maintaining Phase 1 representations or weights during Phase 2, as these incur significant memory overhead and may violate practical data-sharing constraints.